# CSWin-UNet: Transformer UNet with Cross-Shaped Windows for Medical Image Segmentation

### v1

@@ Abstract
Deep learning, especially convolutional neural networks (CNNs) and Transformer architectures, have become the focus of extensive research in medical image segmentation, achieving impressive results. However, CNNs come with inductive biases that limit their effectiveness in more complex, varied segmentation scenarios. Conversely, while Transformer-based methods excel at capturing global and long-range semantic details, they suffer from high computational demands. In this study, we propose CSWin-UNet, a novel U-shaped segmentation method that incorporates the CSWin self-attention mechanism into the UNet to facilitate horizontal and vertical stripes self-attention. This method significantly enhances both computational efficiency and receptive field interactions. Additionally, our innovative decoder utilizes a content-aware reassembly operator that strategically reassembles features, guided by predicted kernels, for precise image resolution restoration. Our extensive empirical evaluations on diverse datasets, including synapse multi-organ CT, cardiac MRI, and skin lesions demonstrate that CSWin-UNet maintains low model complexity while delivering high segmentation accuracy.
Index Terms:
Medical image segmentation, deep learning, attention mechanism, neural network.
I Introduction
@@
The structure of this paper is organized as follows. Section II reviews recent works and developments in the field of medical image segmentation, setting the context for the innovations introduced in this study. Section III describes in detail the methodology of the newly proposed CSWin-UNet, highlighting the novel aspects of the architecture and its components. Section IV presents the experimental results, demonstrating the effectiveness and efficiency of CSWin-UNet compared to existing methods. Section V concludes the paper.
II Related works
II-A Self-attention mechanisms in image segmentation
@@
II-B CNN-based medical image segmentation
@@
II-C Transformer-based medical image segmentation
@@ III Methodology
III-A CSWin Transformer block
@@
$&\text{h}_{n}=\left\{\begin{aligned} \text{H-Attention}_{n}(X)&,\quad n=1,2,\ldots,N/2\\ \text{V-Attention}_{n}(X)&,\quad n=N/2+1,N/2+2,\ldots,N\end{aligned}\right.\\ &\text{CSWin-Attention}(X)=\mathrm{concat}(\text{h}_{1},\text{h}_{2},\ldots,\text{h}_{N})W^{o}\end{aligned}$
@@ III-C Decoder
IV Experiments
IV-A Implementation details
@@
IV-B Datasets and metrics
TABLE I: Detailed comparisons with recent medical image segmentation methods in terms of DSC and HD on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-B Datasets and metrics
IV-B1 Synapse dataset
IV-B2 ACDC dataset
IV-B3 Skin lesion segmentation datasets
IV-C Results on Synapse dataset
@@
As shown in Table I, our proposed method on the Synapse dataset has improved the mean DSC and HD. Meanwhile, we represent error bars (95% confidence interval) of mean DSC, mean HD, and DSC for each organ in Fig.5. Compared to TransUNet[13] and Swin-UNet[30], our mean DSC has increased by 3.64% and 1.99%, respectively, and the mean HD has improved by 12.83% and 2.69%. Notably, in the segmentation of the pancreas, the DSC of CSWin-UNet is significantly higher than that of other segmentation methods. Unlike other organs, the pancreas has blurry boundaries and variability, and our method achieved more precise segmentation results for the pancreas, indicating that our CSWin-UNet provides higher segmentation accuracy in complex segmentation environments.
@@
IV-D Results on ACDC dataset
TABLE II: Comparison results of top-tier medical image segmentation methods on the ACDC dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-D Results on ACDC dataset
Table II presents the experimental results of our proposed CSWin-UNet on the ACDC dataset and compares them with other advanced methods. Fig.7 represents the error bars (95% confidence interval) of mean DSC and DSC for each cardiac structure. In the table, RV represents the right ventricle, MYO represents the myocardium, and LV represents the left ventricle. The results indicate that the proposed CSWin-UNet achieves better identification and segmentation of these organs, with an accuracy rate of 91.40%, demonstrating good generalization capabilities and robustness.
IV-E Results on skin lesion segmentation datasets
TABLE III: Comparative results of the proposed CSWin-UNet with other state-of-the-art methods on three skin lesion segmentation datasets. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-E Results on skin lesion segmentation datasets
Table III shows the experimental results, and Fig.8 shows the error bars (95% confidence interval) of DSC, SE, SP, and ACC on three skin lesion segmentation datasets. The experimental results indicate that the proposed CSWin-UNet outperforms other methods in most evaluation metrics. Notably, CSWin-UNet achieves better performance than Swin-UNet[30] in most metrics, demonstrating satisfactory generalization capability. We also visualized the skin lesion segmentation results in Fig.9. Compared to Swin-UNet[30], our CSWin-UNet has certain advantages in preserving the edges and detailed features of the segmented objects. However, in cases of low contrast or occlusion, as shown in Fig.9(d), the segmentation produces significant errors.
IV-F Comparison of computational efficiency
TABLE IV: Computational efficiency of different medical image segmentation methods on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-F Comparison of computational efficiency
An essential objective for neural network model design is to reduce the parameter count and computational complexity as much as possible while maintaining its performance. This reduction is crucial for enabling more efficient model training and deployment in devices with limited computational resources. Therefore, in evaluating a model, it is necessary to consider not only its accuracy and generalizability but also its parameter count and computational complexity. Here, we use FLOPs and parameters (in millions, M) to measure computational complexity. The performance comparison on the Synapse dataset is shown in Table IV. The results demonstrate that the proposed CSWin-UNet achieves excellent segmentation performance under conditions of the lowest complexity.
IV-G Ablation studies
@@
IV-G1 Upsampling strategy
TABLE V: Ablation study of different upsampling strategies on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-G1 Upsampling strategy
In the encoder, downsampling is performed using a convolutional layer with a stride of 2, and correspondingly, upsampling is needed in the decoder to upsample the feature maps, thereby preserving more information. In this paper, we introduce the CARAFE layer to achieve upsampling and increase feature channels, which uses the content of the input features themselves to guide the upsampling process for more accurate and efficient feature reassembly. To verify the effectiveness of the CAREFE layer, we conducted experiments on the Synapse dataset with bilinear interpolation, transposed convolution, and CARAFE layer in CSWin-UNet, as shown in Table V. Upsampling with the CAREFE layer achieved the highest segmentation accuracy. Moreover, compared to transposed convolution, CAREFE introduces very little computational overhead. Experimental results indicate that the CSWin-UNet combined with the CARAFE layer can achieve optimal performance.
IV-G2 Skip connection
TABLE VI: Ablation study of different numbers of skip connection on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-G2 Skip connection
Similar to the UNet, we also introduced skip connections to enhance finer segmentation details by restoring low-level spatial information. In CSWin-UNet, skip connections are positioned at the resolution scales of 1/4, 1/8, and 1/16. We sequentially reduced the skip connections at the scales of 1/16, 1/8, and 1/4, setting the number of skip connections to 3, 2, 1, and 0 to explore the impact of varying numbers of skip connections on segmentation accuracy. As shown in Table VI, segmentation accuracy generally improves with an increase in the number of skip connections. Notably, CSWin-UNet achieved more significant improvements in segmentation accuracy on smaller organs (such as the aorta, gallbladder, kidneys, and pancreas) than on larger organs (such as the liver, spleen, and stomach). Therefore, to achieve optimal performance, we set the number of skip connections to 3.
IV-G3 Network architecture
TABLE VII: Ablation study of different numbers of CSWin Transformer blocks of each stage on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-G3 Network architecture
Neural networks with too few layers can result in insufficiently rich and accurate feature representations, making it difficult to understand the image context, thereby leading to poor segmentation performance. Conversely, having too many layers increases the computational burden and makes it difficult for the network to converge. Therefore, when designing the network architecture, a balance was struck between network depth and model performance, enabling the model to achieve high segmentation accuracy with limited computational resources. Additionally, to prevent non-convergence issues due to excessive depth[55], the block count in the final stage is set to one. By comparing the number of parameters and computational costs of other Transformer-based medical image segmentation methods, we set the block numbers in the four stages to [1, 2, 6, 1], [1, 2, 9, 1], and [1, 2, 12, 1], with encoder and decoder blocks symmetrically arranged. As shown in Table VII, the network architecture with block settings of [1, 2, 9, 1] achieved the best performance.
IV-G4 Combined loss function
TABLE VIII: Ablation study of different hyperparameters of the combined loss function on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ IV-G4 Combined loss function
We explored the impact of different hyperparameters of the combined loss function on segmentation accuracy. Here, we set $\alpha$ and $\beta$ in Eq.10 to $[1,0]$, $[0,1]$, $[0.5,0.5]$, $[0.4,0.6]$, and $[0.3,0.7]$. We conducted an ablation study on the Synapse dataset, and the experimental results showed that using the combined loss function resulted in higher segmentation accuracy than using either Dice or cross-entropy losses alone, especially in the case of using only Dice loss without cross-entropy loss. Table VIII shows that the segmentation performance was optimal when $\alpha$ and $\beta$ were set to $[0.4,0.6]$.
IV-H Discussions

### v3

@@ Abstract
Deep learning, especially convolutional neural networks (CNNs) and Transformer architectures, have become the focus of extensive research in medical image segmentation, achieving impressive results. However, CNNs come with inductive biases that limit their effectiveness in more complex, varied segmentation scenarios. Conversely, while Transformer-based methods excel at capturing global and long-range semantic details, they suffer from high computational demands. In this study, we propose CSWin-UNet, a novel U-shaped segmentation method that incorporates the CSWin self-attention mechanism into the UNet to facilitate horizontal and vertical stripes self-attention. This method significantly enhances both computational efficiency and receptive field interactions. Additionally, our innovative decoder utilizes a content-aware reassembly operator that strategically reassembles features, guided by predicted kernels, for precise image resolution restoration. Our extensive empirical evaluations on diverse datasets, including synapse multi-organ CT, cardiac MRI, and skin lesions, demonstrate that CSWin-UNet maintains low model complexity while delivering high segmentation accuracy. Codes are available at https://github.com/eatbeanss/CSWin-UNet.
1 Introduction
@@
The structure of this paper is organized as follows. Section 2 reviews recent works and developments in the field of medical image segmentation, setting the context for the innovations introduced in this study. Section 3 describes in detail the methodology of the newly proposed CSWin-UNet, highlighting the novel aspects of the architecture and its components. Section 4 presents the experimental results, demonstrating the effectiveness and efficiency of CSWin-UNet compared to existing methods. Section 5 concludes the paper.
2 Related works
2.1 Self-attention mechanisms in image segmentation
@@
2.2 CNN-based medical image segmentation
@@
2.3 Transformer-based medical image segmentation
@@ 3 Methodology
3.1 CSWin Transformer block
@@
$\text{h}_{n}=\left\{\begin{aligned} \text{H-Attention}_{n}(X)&,\quad n=1,2,\ldots,N/2\\ \text{V-Attention}_{n}(X)&,\quad n=N/2+1,N/2+2,\ldots,N\end{aligned}\right.$ (5)
[…]
@@ 3.3 Decoder
4 Experiments
4.1 Implementation details
@@
4.2 Datasets and metrics
4.2.1 Synapse dataset
4.2.2 ACDC dataset
4.2.3 Skin lesion segmentation datasets
4.3 Results on Synapse dataset
Table 1: Detailed comparisons with recent medical image segmentation methods in terms of DSC and HD on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.3 Results on Synapse dataset
As shown in Table 1, our proposed method on the Synapse dataset has improved the mean DSC and HD. Meanwhile, we represent error bars (95% confidence interval) of mean DSC, mean HD, and DSC for each organ in Fig.5. Compared to TransUNet[13] and Swin-UNet[30], our mean DSC has increased by 3.64% and 1.99%, respectively, and the mean HD has improved by 12.83% and 2.69%. Notably, in the segmentation of the pancreas, the DSC of CSWin-UNet is significantly higher than that of other segmentation methods. Unlike other organs, the pancreas has blurry boundaries and variability, and our method achieved more precise segmentation results for the pancreas, indicating that our CSWin-UNet provides higher segmentation accuracy in complex segmentation environments.
@@
4.4 Results on ACDC dataset
Table 2: Comparison results of top-tier medical image segmentation methods on the ACDC dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.4 Results on ACDC dataset
Table 2 presents the experimental results of our proposed CSWin-UNet on the ACDC dataset and compares them with other advanced methods. Fig.7 represents the error bars (95% confidence interval) of mean DSC and DSC for each cardiac structure. In the table, RV represents the right ventricle, MYO represents the myocardium, and LV represents the left ventricle. The results indicate that the proposed CSWin-UNet achieves better identification and segmentation of these organs, with an accuracy rate of 91.40%, demonstrating good generalization capabilities and robustness.
4.5 Results on skin lesion segmentation datasets
Table 3: Comparative results of the proposed CSWin-UNet with other state-of-the-art methods on three skin lesion segmentation datasets. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.5 Results on skin lesion segmentation datasets
Table 3 shows the experimental results, and Fig.8 shows the error bars (95% confidence interval) of DSC, SE, SP, and ACC on three skin lesion segmentation datasets. The experimental results indicate that the proposed CSWin-UNet outperforms other methods in most evaluation metrics. Notably, CSWin-UNet achieves better performance than Swin-UNet[30] in most metrics, demonstrating satisfactory generalization capability. We also visualized the skin lesion segmentation results in Fig.9. Compared to Swin-UNet[30], our CSWin-UNet has certain advantages in preserving the edges and detailed features of the segmented objects. However, in cases of low contrast or occlusion, as shown in Fig.9(d), the segmentation produces significant errors.
4.6 Comparison of computational efficiency
Table 4: Computational efficiency of different medical image segmentation methods on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.6 Comparison of computational efficiency
An essential objective for neural network model design is to reduce the parameter count and computational complexity as much as possible while maintaining its performance. This reduction is crucial for enabling more efficient model training and deployment in devices with limited computational resources. Therefore, in evaluating a model, it is necessary to consider not only its accuracy and generalizability but also its parameter count and computational complexity. Here, we use FLOPs and parameters (in millions, M) to measure computational complexity. The performance comparison on the Synapse dataset is shown in Table 4. The results demonstrate that the proposed CSWin-UNet achieves excellent segmentation performance under conditions of the lowest complexity.
4.7 Ablation studies
@@
4.7.1 Upsampling strategy
Table 5: Ablation study of different upsampling strategies on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.7.1 Upsampling strategy
In the encoder, downsampling is performed using a convolutional layer with a stride of 2, and correspondingly, upsampling is needed in the decoder to upsample the feature maps, thereby preserving more information. In this paper, we introduce the CARAFE layer to achieve upsampling and increase feature channels, which uses the content of the input features themselves to guide the upsampling process for more accurate and efficient feature reassembly. To verify the effectiveness of the CAREFE layer, we conducted experiments on the Synapse dataset with bilinear interpolation, transposed convolution, and CARAFE layer in CSWin-UNet, as shown in Table 5. Upsampling with the CAREFE layer achieved the highest segmentation accuracy. Moreover, compared to transposed convolution, CAREFE introduces very little computational overhead. Experimental results indicate that the CSWin-UNet combined with the CARAFE layer can achieve optimal performance.
4.7.2 Skip connection
Table 6: Ablation study of different numbers of skip connection on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.7.2 Skip connection
Similar to the UNet, we also introduced skip connections to enhance finer segmentation details by restoring low-level spatial information. In CSWin-UNet, skip connections are positioned at the resolution scales of 1/4, 1/8, and 1/16. We sequentially reduced the skip connections at the scales of 1/16, 1/8, and 1/4, setting the number of skip connections to 3, 2, 1, and 0 to explore the impact of varying numbers of skip connections on segmentation accuracy. As shown in Table 6, segmentation accuracy generally improves with an increase in the number of skip connections. Notably, CSWin-UNet achieved more significant improvements in segmentation accuracy on smaller organs (such as the aorta, gallbladder, kidneys, and pancreas) than on larger organs (such as the liver, spleen, and stomach). Therefore, to achieve optimal performance, we set the number of skip connections to 3.
4.7.3 Network architecture
Table 7: Ablation study of different numbers of CSWin Transformer blocks of each stage on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.7.3 Network architecture
Neural networks with too few layers can result in insufficiently rich and accurate feature representations, making it difficult to understand the image context, thereby leading to poor segmentation performance. Conversely, having too many layers increases the computational burden and makes it difficult for the network to converge. Therefore, when designing the network architecture, a balance was struck between network depth and model performance, enabling the model to achieve high segmentation accuracy with limited computational resources. Additionally, to prevent non-convergence issues due to excessive depth[55], the block count in the final stage is set to one. By comparing the number of parameters and computational costs of other Transformer-based medical image segmentation methods, we set the block numbers in the four stages to [1, 2, 6, 1], [1, 2, 9, 1], and [1, 2, 12, 1], with encoder and decoder blocks symmetrically arranged. As shown in Table 7, the network architecture with block settings of [1, 2, 9, 1] achieved the best performance.
4.7.4 Combined loss function
Table 8: Ablation study of different hyperparameters of the combined loss function on the Synapse dataset. The first and second best values are highlighted in red and blue fonts, respectively.
@@ 4.7.4 Combined loss function
We explored the impact of different hyperparameters of the combined loss function on segmentation accuracy. Here, we set $\alpha$ and $\beta$ in Eq.10 to $[1,0]$, $[0,1]$, $[0.5,0.5]$, $[0.4,0.6]$, and $[0.3,0.7]$. We conducted an ablation study on the Synapse dataset, and the experimental results showed that using the combined loss function resulted in higher segmentation accuracy than using either Dice or cross-entropy losses alone, especially in the case of using only Dice loss without cross-entropy loss. Table 8 shows that the segmentation performance was optimal when $\alpha$ and $\beta$ were set to $[0.4,0.6]$.
4.8 Discussions